OpenStudy (anonymous):

Use DeMoivre's Theorem to write the complex number in trigonometric form. (sqrt2 + I sqrt2)^3

OpenStudy (anonymous):

@ zepdrix

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

So we need 2 pieces of information: The magnitude, \(\Large\rm r\), of our thing inside the brackets, and the angle it forms \(\Large\rm \theta\).

zepdrix (zepdrix):

So when we have the form:\[\Large\rm a+b\mathcal i\](Which we do) Then the radius is given by: \[\Large\rm r=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

so that would be 2

zepdrix (zepdrix):

Mmm ok good.

OpenStudy (anonymous):

r=2

zepdrix (zepdrix):

Recall that tangent is \(\Large\rm \dfrac{opposite}{adjacent}\) What that means for us here is that the tangent of theta is the imaginary component (opposite side) over the real component (adjacent side). \[\Large\rm \tan \theta=\frac{b}{a}\]

zepdrix (zepdrix):

Website crashed again? :( ugh great...

OpenStudy (anonymous):

that equals 1

OpenStudy (anonymous):

@zepdrix you still there

zepdrix (zepdrix):

\[\Large\rm \tan \theta=1\]Yes, good. So how do we solve for theta? What special angles does this correspond to?